The ratio of Anne's height :
Ben's height is 7 : 9.
Anne's height is 1.4 metres.
Find Ben's height.

Answer by geetha_rama(94)

Let Ben's height by y meters
Ratio of Annie height and ben height is 7:9
hence 7:9 :: 1.4:y
=>7/9 = 1.4/y
=> 7y = 1.4 * 9
=>y = 1.4* 9/7
= .2*9
=> y = 1.8
hence Ben's Height is 1.8 metres